Saturn and Jupiter take 9 hours 56 minutes and 10 hours 40 minutes, respectively for one spin on their axes. The ratio of the time taken by Saturn and Jupiter in lowest form is ______.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Saturn and Jupiter take 9 hours 56 minutes and 10 hours 40 minutes, respectively for one spin on their axes. The ratio of the time taken by Saturn and Jupiter in lowest form is 149: 160.
From the question,

Saturn takes 9 hours 56 minutes for one spin on their axes We know that, 1 hour = 60 minutes
So, (9 × 60) + 56 = 540 + 56 = 596 minutes
Jupiter takes 10 hours 40 minutes for one spin on their axes

= (10 × 60) + 40 = 600 + 40
= 640 minutes

The ratio of the time taken by Saturn and Jupiter in lowest form is = 596/640 Divide both numerator and denominator by 2,

= 298/320
Again, divide both numerator and denominator by 2,

= 149/160
Therefore, the ratio of the time taken by Saturn and Jupiter in lowest form is 149: 160.

Thanks